Is matrix A consistent, motivate your answer in terms of eigenvalues. A = 2 1 0 0 2 0 2 3 1
1
Expert's answer
2021-10-26T08:55:22-0400

If "A" is an "n \u00d7 n" matrix, then the sum of the "n" eigenvalues of "A" is the trace of "A" and the product of the "n" eigenvalues is the determinant of "A."


"A=\begin{pmatrix}
 2 & 1 & 0 \\
 0 & 2 & 0\\
 2 &3 & 1
\end{pmatrix}"

"A-\lambda I=\begin{pmatrix}
 2-\lambda & 1 & 0 \\
 0 & 2-\lambda & 0\\
 2 &3 & 1-\lambda
\end{pmatrix}"

"\det (A-\lambda I)=\begin{vmatrix}
 2-\lambda & 1 & 0 \\
 0 & 2-\lambda & 0\\
 2 &3 & 1-\lambda
\end{vmatrix}"

"=(2-\lambda)\begin{vmatrix}
 2-\lambda & 0 \\
 3 & 1-\lambda
\end{vmatrix}-1\begin{vmatrix}
 0 & 0 \\
 2 & 1-\lambda
\end{vmatrix}+0\begin{vmatrix}
 0 & 2-\lambda \\
 2 & 3
\end{vmatrix}"

"=(2-\lambda)(2-\lambda)(1-\lambda)"

Characteristic equation


"\det (A-\lambda I)=0"

"(2-\lambda)(2-\lambda)(1-\lambda)=0"

"\lambda_1=1, \lambda_2=2, \lambda_3=2"

These are the eigenvalues.

Hence


"\det A=\lambda_1\lambda_2\lambda_3=1(2)(2)=4\not=0"

=> matrix "A" consistent.